An aerobatic airplane pilot experiences
weightlessness as she passes over the top of
a loop-the-loop maneuver.
The acceleration of gravity is 9.8 m/s2. If her speed is 310 m/s at this time, find the
radius of the loop.
Answer in units of km.
1
Expert's answer
2019-12-23T11:42:27-0500

There are two forces that act on the pilot: the gravitational force "mg" directed towards the center of the loop and the normal force exerted on the pilot "N" directed away from the center of the loop. Let's choose the downwards as the positive direction and apply the Newton's Second Law of Motion:


"mg - N = m \dfrac{v^2}{r}."

Since the pilot experiences the weightlessness at the top point of the loop, the normal force exerted on the pilot equals zero. Therefore, we get:


"g = \dfrac{v^2}{r}."

From this formula we can find the radius of the loop:

"r = \dfrac{v^2}{g} = \dfrac{(310 \dfrac{m}{s})^2}{9.8 \dfrac{m}{s^2}} = 9.806 \cdot 10^3 m = 9.8 km."

Answer:

"r = 9.8 km."